at the beginning of the summer, sarah has $250. she takes a summer job and saves $150 per week. felicia has $1,650 at the beginning of the summer. she travels during the summer and spends $200 per week. write an equation that represents the amount of money sarah and felicia has at the end of each week.
Answer by Theo(13342)

at the end of each week, sarah has:

250 + (150*w)

at the end of each week, felicia has:

1650 - (200*w)

w = number of weeks.

at the end of the first week:

sarah has 250 + 1*150 = 400
felicia has 1650 - 1*200 = 1450

at the end of the second week:

sarah has 250 + 2*150 = 550
felicia has 1650 - 2*200 = 1250

etc.